### v1

@@ Abstract
Due to their importance in various emerging applications, efficient algorithms for solving minimax problems have recently received increasing attention. However, many existing algorithms require prior knowledge of the problem parameters in order to achieve optimal iteration complexity. In this paper, we propose a completely parameter-free alternating gradient projection (PF-AGP) algorithm to solve the smooth nonconvex-(strongly) concave minimax problems using a backtracking strategy, which does not require prior knowledge of parameters such as the Lipschtiz constant $L$ or the strongly concave constant $\mu$. The PF-AGP algorithm utilizes a parameter-free gradient projection step to alternately update the outer and inner variables in each iteration. We show that the total number of gradient calls of the PF-AGP algorithm to obtain an $\varepsilon$-stationary point for nonconvex-strongly concave minimax problems is upper bounded by $\mathcal{O}\left(L\kappa^{3}\varepsilon^{-2}\right)$ where $\kappa$ is the condition number, while the total number of gradient calls to obtain an $\varepsilon$-stationary point for nonconvex-concave minimax problems is upper bounded by $\mathcal{O}\left(L^{4}\varepsilon^{-4}\right)$. As far as we know, this is the first completely parameter-free algorithm for solving nonconvex-strongly concave minimax problems, and it is also the completely parameter-free algorithm which achieves the best iteration complexity in single loop method for solving nonconvex-concave minimax problems. Numerical results validate the efficiency of the proposed PF-AGP algorithm.
@@ 1 Introduction
where $\mathcal{X}\subseteq\mathbb{R}^{n}$ and $\mathcal{Y}\subseteq\mathbb{R}^{m}$ are nonempty closed and bounded convex sets, and $f:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ is a smooth function. The minimax problem is a critical class of optimization problems, known for its extensive research history and broad applications in areas such as machine learning, signal processing, and various other fields in recent years. Many practical problems can be formulated as in (P), including machine learning, such as generative adversarial networks (GANs) (Arjovsky et al., 2017; Goodfellow et al., 2014; Sanjabi et al., 2018), reinforcement learning (Dai et al., 2018), power control and transceiver design in signal processing (Lu et al., 2020), distributed nonconvex optimization (Giannakis et al., 2017; Giordano et al., 2018; Mateos et al., 2010), robust optimization (Ben and Tal, 2009; Gao and Kleywegt, 2023), statistical learning (Giordano et al., 2018; Shafieezadeh Abadeh et al., 2015) and several other applications.
@@ 1.2 Organization
In Sect. 2, we propose a parameter-free alternating gradient projection (PF-AGP) algorithm for nonconvex-(strongly) concave minimax problems, and we then analyze the corresponding gradient complexity for two different settings. We report some numerical results in Sect. 3 and make some concluding remarks in the last section.
@@ 2.1 Nonconvex-strongly Concave Setting
It is worth noting that the parameter-free NeAda (Yang et al., 2022) and TiAda (Li et al., 2022) algorithms with adaptive step size are only used to solve the minimax problems when $x$ is unconstrained, while the proposed PF-AGP-NSC algorithm can solve the constrained case. Compared with the SGDA-B (Xu et al., 2023) algorithm and AGDA+ (Zhang et al., 2024) algorithm using the backtracking framework under the nonconvex-strongly concave setting, AGDA+ and SGDA-B still require knowledge of the strongly concave constant $\ mu$, while the proposed PF-AGP-NSC algorithm is a completely parameter-free algorithm that does not require any prior knowledge of the parameters.
@@ 2.1.1 Complexity analysis
In this subsection, we analyze the iteration complexity of Algorithm 1 for solving nonconvex-strongly concave minimax optimization problems, i.e., $f(x,y)$ is nonconvex with respect to $x$ for any fixed $y\in\mathcal{Y}$, and $\mu$-strongly concave with respect to $y$ for any given $x\in\mathcal{X}$.
Before we prove the iteration complexity of the PF-AGP-NSC algorithm for solving (P), we need to make the following assumption about the smoothness of $f(x,y)$.
@@ 3 Numerical results
In this section, we compare the numerical performance of the proposed PF-AGP-NC algorithm with the AGP algorithm.
@@ 3.1 Dirac-GAN problem
The Dirac-GAN problem (Mescheder et al., 2018) can be formulated as the following nonconvex–concave minimax problem :
@@
Experimental setup. We set the initial point of the experiment at $(x,y)=(1,1)$. The termination condition is defined as the norm of the gradient $\|\nabla f(x_{i},y_{i})\|$ being less than or equal to $10^{-5}$, where $i$ represents the number of gradient calls.
For the AGP algorithm, let $\alpha_{x}$ and $\beta_{y}$ be the stepsize of $x$ and $y$, respectively. We set $\alpha_{x}=\frac{0.8}{\sqrt{k}}$, $\beta_{y}=0.3$ and $c_{k}=\frac{0.5}{k^{1/4}}$. For the PF-AGP-NC algorithm, we do not need to preset specific parameters.
In the experiments, we counted the number of gradient calls of the PF-AGP-NC algorithm. The results of alternating gradient projection in the $k$ th iteration are represented by nodes in the graph and the number of backtracking step computations is represented by the connecting lines.
Results Figure 1 shows the convergence curves of the AGP and PF-AGP-NC algorithms. It can be seen that although the PF-AGP-NC algorithm involves a double-layer structure, requiring the determination of points in the outer layer through computations in the inner layer, even when considering the number of iterations in the inner layer, the overall convergence speed is still faster than that of the AGP algorithm.
Figure 1: number of gradient calls $i$ vs $\log(||\nabla f(x_{i},y_{i})||)$
4 Conclusions and discussion

### v2

@@ Abstract
Due to their importance in various emerging applications, efficient algorithms for solving minimax problems have recently received increasing attention. However, many existing algorithms require prior knowledge of the problem parameters in order to achieve optimal iteration complexity. In this paper, we propose two completely parameter-free alternating gradient projection algorithms, i.e., the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm, to solve the smooth nonconvex-strongly concave and nonconvex-concave minimax problems respectively using a backtracking strategy, which does not require prior knowledge of parameters such as the Lipschtiz constant $L$ or the strongly concave constant $\mu$. Moreover, we show that the total number of gradient calls of the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm to obtain an $\varepsilon$-stationary point is upper bounded by $\mathcal{O}\left(L\kappa^{3}\varepsilon^{-2}\right)$ and $\mathcal{O}\left(L^{4}\varepsilon^{-4}\right)$ respectively, where $\kappa$ is the condition number. As far as we know, the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm are the first completely parameter-free algorithms for solving nonconvex-strongly concave minimax problems and nonconvex-concave minimax problems respectively. Numerical results validate the efficiency of the proposed PF-AGP algorithm.
@@ 1 Introduction
where $\mathcal{X}\subseteq\mathbb{R}^{n}$ and $\mathcal{Y}\subseteq\mathbb{R}^{m}$ are nonempty closed and bounded convex sets, and $f:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ is a smooth function. Minimax problem is a critical class of optimization problems, known for its extensive research history and broad applications in areas such as machine learning, signal processing, and various other fields in recent years. Many practical problems can be formulated as in (P), including generative adversarial networks (GANs) (Arjovsky et al., 2017; Goodfellow et al., 2014; Sanjabi et al., 2018), reinforcement learning (Dai et al., 2018), power control and transceiver design in signal processing (Lu et al., 2020), distributed nonconvex optimization (Giannakis et al., 2017; Giordano et al., 2018; Mateos et al., 2010), robust optimization (Ben and Tal, 2009; Gao and Kleywegt, 2023), statistical learning (Giordano et al., 2018; Shafieezadeh Abadeh et al., 2015) and several other applications.
@@ 1.2 Organization
In Sect. 2, we propose two parameter-free alternating gradient projection (PF-AGP) algorithms for nonconvex-strongly concave and nonconvex-concave minimax problems, and we then analyze their corresponding gradient complexity under two different settings. We report some numerical results in Sect. 3 and make some conclusions in the last section.
@@ 2.1 Nonconvex-strongly Concave Setting
It is worth noting that the parameter-free NeAda (Yang et al., 2022) and TiAda (Li et al., 2022) algorithms with adaptive step size are only used to solve the minimax problems when $x$ is unconstrained, while the proposed PF-AGP-NSC algorithm can solve the constrained case. Compared with the SGDA-B (Xu et al., 2023) algorithm and AGDA+ (Zhang et al., 2024) algorithm using the backtracking framework under the nonconvex-strongly concave setting, AGDA+ and SGDA-B still require knowledge of the strongly concave constant $\mu$, while the proposed PF-AGP-NSC algorithm is a completely parameter-free algorithm that does not require any prior knowledge of the parameters.
@@ 2.1.1 Complexity analysis
In this subsection, we analyze the iteration complexity of Algorithm 1 for solving nonconvex-strongly concave minimax optimization problems (P), i.e., $f(x,y)$ is nonconvex with respect to $x$ for any fixed $y\in\mathcal{Y}$, and $\mu$-strongly concave with respect to $y$ for any given $x\in\mathcal{X}$. We first need to make the following assumption about the smoothness of $f(x,y)$.
@@ 3 Numerical results
In this section, we report some numerical results to show the efficiency of the proposed PF-AGP-NSC and PF-AGP-NC algorithms. All the numerical tests are implemented in Python 3.12 and run with an NVIDIA RTX 4090D GPU (24GB).
3.1 Synthetic Minimax Problem
Synthetic minimax problem can be formulated as the following nonconvex-strongly-concave minimax problem (Luo et al., 2022):
$\min_{\mathbf{x}\in\mathbb{R}^{3}}\max_{\mathbf{y}\in\mathbb{R}^{2}}f(\mathbf{x},\mathbf{y})=w(x_{3})-\frac{y_{1}^{2}}{40}+x_{1}y_{1}-\frac{5y_{2}^{2}}{2}+x_{2}y_{2},$ (82)
where $\mathbf{x}=[x_{1},x_{2},x_{3}]^{\top}$, $\mathbf{y}=[y_{1},y_{2}]^{\top}$ and
$w(x)=\begin{cases}\sqrt{\epsilon}\left(x+(\lambda+1)\sqrt{\epsilon}\right)^{2}-\frac{1}{3}\left(x+(\lambda+1)\sqrt{\epsilon}\right)^{3}-\frac{1}{3}(3\lambda+1)\epsilon^{3/2},&x\leq-\lambda\sqrt{\epsilon};\\ \epsilon x+\frac{\epsilon^{3/2}}{3},&-\lambda\sqrt{\epsilon}<x\leq-\sqrt{\epsilon};\\ -\sqrt{\epsilon}x^{2}-\frac{x^{3}}{3},&-\sqrt{\epsilon}<x\leq 0;\\ -\sqrt{\epsilon}x^{2}+\frac{x^{3}}{3},&0<x\leq\sqrt{\epsilon};\\ -\epsilon x+\frac{\epsilon^{3/2}}{3},&\sqrt{\epsilon}<x\leq\lambda\sqrt{\epsilon};\\ \sqrt{\epsilon}\left(x-(\lambda+1)\sqrt{\epsilon}\right)^{2}+\frac{1}{3}\left(x-(\lambda+1)\sqrt{\epsilon}\right)^{3}-\frac{1}{3}(3\lambda+1)\epsilon^{3/2},&\lambda\sqrt{\epsilon}\leq x.\end{cases}$
It is easy to verify that this problem has a strict saddle point at $(\mathbf{x}^{*},\mathbf{y}^{*})=([0,0,0]^{\top},[0,0]^{\top})$.
Experimental setup. We compare the proposed PF-AGP-NSC algorithm with the SGDA-B algorithm, the TiAda algorithm and the AGP algorithm for solving (82) with $\epsilon=0.01$, $\lambda=5$. The initial point is chosen as $(x_{0},y_{0})=([0,0,2]^{\top},[0,0]^{\top})$ for all four tested algorithms. For the PF-AGP-NSC algorithm, we set $l_{11}^{1}=0.01$ , $l_{12}^{1}=0.01$, $l_{22}^{1}=0.01$ and $\mu_{1}=1$. For the SGDA-B algorithm, we set $\mu=1$, $\gamma=0.9$, $N=1$, $p=0.5$ and $\tilde{L}=\frac{10}{9}$. For the TiAda algorithm, we set $\alpha=0.6$, $\beta=0.4$, $\eta_{x}=0.8$, $\eta_{y}=0.4$, $v^{x}_{0}=v^{y}_{0}=1$. For the AGP algorithm, we set $\alpha_{x}=0.14$, $\beta_{y}=1.1$. The algorithms terminate when $\|\nabla f(x_{k},y_{k})\|\leq 10^{-5}$ satisfied.
Figure 1: Numerical Performance of four tested algorithms for solving the Synthetic Minimax problem
Results. Figure 1 shows the stationarity gap of all four test algorithms, which demonstrates that the proposed PF-AGP-NSC algorithm exhibits significantly better convergence performance than that of the SGDA-B algorithm, TiAda algorithm and AGP algorithm.
@@ 3.2 Dirac-GAN problem
The Dirac-GAN problem can be formulated as the following nonconvex-concave minimax problem (Mescheder et al., 2018):
@@
Experimental setup. We compare the PF-AGP-NC algorithm with the SGDA-B algorithm and the AGP algorithm. The initial point of all algorithms is chosen as $(1,1)$. For the PF-AGP-NC algorithm, we set $l_{11}^{1}=0.01$ , $l_{12}^{1}=1$ and $l_{22}^{1}=0.01$. For the SGDA-B algorithm, we set $\mu=1$, $\gamma=0.5$, $N=1$, $p=0.5$ and $\tilde{L}$ = 1. For the AGP algorithm, we set $\alpha_{x}=\frac{0.8}{\sqrt{k}}$, $\beta_{y}=0.3$ and $c_{k}=\frac{0.5}{k^{1/4}}$. The algorithms terminate when $\|\nabla f(x_{k},y_{k})\|\leq 10^{-5}$ satisfied.
Results. Figure 2 shows the stationarity gap and the sequences generated by all three test algorithms. We find that the SGDA-B algorithm fails to converge, both the AGP algorithm and the PF-AGP-NC algorithms successfully approach the unique stationary point. Moreover, the PF-AGP-NC algorithm exhibits significantly better convergence performance than that of the AGP algorithm.
Figure 2: Numerical performance of three tested algorithms for solving the Dirac-GAN problem
3.3 Robust learning over multiple domains
In this subsection, we consider the following robust learning problem over multiple domains (Qian et al., 2019), formulated as a nonconvex-linear minimax problem:
$\min_{x}\max_{y\in\Delta}y^{T}F(x),$ (84)
where $F(x):=\left[f_{1}(x);\ldots;f_{M}(x)\right]\in\mathbb{R}^{M\times 1}$ with $f_{m}(x)=\frac{1}{\left|\mathcal{S}_{m}\right|}\sum_{i=1}^{\left|\mathcal{S}_{m}\right|}\ell\left(d_{i}^{m},l_{i}^{m},x\right)$ and $\ell$ can be any non-negative loss function, $M$ is the number of tasks and $x$ denotes the network parameters, $y\in\Delta$ describes the weights over different tasks and $\Delta$ is the simplex, i.e., $\Delta=\left\{(y_{1},\cdots,y_{M})^{T}\mid\sum_{m=1}^{M}y_{m}=1,y_{m}\geq 0\right\}$. We consider two image classification problems with MNIST (LeCun et al., 1998) and CIFAR10 datasets (Krizhevsky et al., 2009). Our goal is to train a neural network that works on these two completely unrelated problems simultaneously. Since cross-entropy loss is popular in multi-class classification problem, we use it as our loss function. We should point out that the quality of the algorithm for solving a robust learning problem over multiple domains is measured by the worst case accuracy over all domains (Qian et al., 2019).
Experiment setup. We compare the proposed PF-AGP-NC algorithm with the SGDA-B algorithm and the AGP algorithm. For the PF-AGP-NC algorithm, we set $l_{11}^{1}=0.01$, $l_{12}^{1}=0.01$ and $l_{22}^{1}=0.01$. For the SGDA-B algorithm, we set $\mu=20$, $\gamma=0.9$, $N=1$, $p=0.5$ and $\tilde{L}$ = 20. For the AGP algorithm, we set $\frac{1}{\beta_{k}}=\frac{2}{2+\sqrt{k}}$, $\gamma_{k}=100$ and $c_{k}=\frac{1}{10+k^{\frac{1}{4}}}$. We set the batch size in all tasks to be $128$ and run $50$ epochs for all algorithms. Moreover, we sample our results every epoch and calculate the average accuracy on these two testing sets to evaluate the performance of different algorithms. Due to hardware (especially memory) limitations, we use a mini-batch randomized gradient instead of the exact gradient at each iteration in our numerical experiment, which is the same as in (Xu et al., 2023).
Figure 3: Performance of two algorithms for solving robust multi-task learning problem.
Results. Figure 3 and Figure 3 show the testing accuracy on the MNIST dataset and the CIFAR10 dataset respectively. PF-AGP-NC algorithm outperforms the SGDA-B algorithm and the AGP algorithm for both datasets.
4 Conclusions